simplify the following 3/5+7/3+(-13/15)+2/5+(-2/3)
Answer by fcabanski(1391)

When adding or subtracting all fractions must have the same denominator. Find the least common denominator, or at least a common denominator.


In this case it's clear the common denominator is 15.


3/5 = 9/15; 7/3 = 35/15; -13/15 stays as it is; 2/5 = 6/15; -2/3 = -10/15. Now add the new fractions:


9/15 + 35/15 + (-13/15) + 6/15 + (-10/15) = 27/15 = 9/5

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)